# Anatomy of Elite and Mass Polarization in Social Networks

### v1

@@ Abstract
Existing methods for quantifying polarization in social networks typically report a single value describing the amount of polarization in a social system. While this approach can be used to confirm the observation that many societies have witnessed an increase in political polarization in recent years, it misses the complexities that could be used to understand the reasons behind this phenomenon. Notably, opposing groups can have unequal impact on polarization, and the elites are often understood to be more divided than the masses, making it critical to differentiate their roles in polarized systems. We propose a method to characterize these distinct hierarchies in polarized networks, enabling separate polarization measurements for these groups within a single social system. Applied to polarized topics in the Finnish Twittersphere surrounding the 2019 and 2023 parliamentary elections, our analysis reveals valuable insights: 1) The impact of opposing groups on observed polarization is rarely balanced, and 2) while the elite strongly contributes to structural polarization and consistently display greater alignment across various topics, the masses have also recently experienced a surge in issue alignment, a special form of polarization. Our findings suggest that the masses may not be as immune to an increasingly polarized environment as previously thought.
@@ 1 Introduction
Distinct aspects of political polarization are reportedly higher today in many societies [ILL+19, Sil18, KYA21, FZQ+23, BG20, KC24]. It is a serious and consequential phenomenon that has garnered significant attention in recent years [FGT+22, XGM+22, FGF+23, Kaw22]. It refers to the increasing divergence of political attitudes, beliefs, and values between individuals and groups [DEB96, FA08], often leading to heightened hostility [FH23, BC18], ideological rigidity [WCY+22], and a breakdown in constructive dialogue [FH23]. As polarization intensifies, it poses significant challenges to democratic societies [MRS18], hindering effective governance [Jon01] and compromising social cohesion [Mas15]. Hence the implications of political polarization are far-reaching, affecting policy-making [WHD+21], public discourse [RM16], and even personal relationships [WCP21, GS23]. Understanding the causes, dynamics, and consequences of political polarization is crucial in order to address its negative impacts.
Current methods for quantifying polarization in social networks assign a single score to an entire social system (e.g., [GMGM18, GMJCK13, MBLB15, LAD24, GAS+15]), simplifying the complexity of the underlying phenomenon. While these methods are often intuitive and straightforward to implement, they produce high polarization values for random networks [SCK22] and fail to capture many crucial nuances related to the multifaceted nature of polarization. They do not distinguish between the elite and the mass polarization, nor do they reveal whether opposing groups contribute equally to the observed division in a given network. These methods are also typically designed to measure the decreasing communication between groups and/or increasing cohesion within them, thus missing other aspects of polarization, such as political alignment [Mas15, Tör22, CSG+21].
We argue that knowing whether both conceptually and structurally dominant individuals are more polarized than the less dominant individuals is important for understanding such polarized systems. Moreover, describing overall polarization with a single score can create the false perception that all groups are equally influential in the observed polarization online. Despite a growing body of work reporting increasing polarization worldwide [AR21, CH23, KYA21, FGT+22, KC24, BG20], we lack standardized methods to break down the measured division into more interpretable components that provide deeper insights into the phenomenon. In this study, we aim to extend the informativity of common polarization measures. We begin by identifying the assortative clusters [ZP20] representing the polarized groups. Then, we further partition each identified group into a hierarchy using the well-known core-periphery structure [BE00, RPFM17, GYW21] to separate the elite from the mass. Finally, we study two distinct aspects of polarization: structural polarization and issue alignment, at both the group and hierarchical levels. For the former, this involves decomposing an existing polarization measure, while for the latter, we measure the degree of alignment separately for the identified hierarchical groups.
Some polarization studies focus solely on the political elite (e.g. [Nea20]), such as elected officials, media figures or influential individuals, while some are more limited to assess the division among the general mass (e.g. [FA08, Lel16]). Researchers have also explored the interplay between these two groups (elite and mass) in polarized settings (e.g. [BC18, BCRR23, Sky21, RM16, FA08, LC02]), resulting in valuable insights into their differing levels of political responsiveness [BCN+19] and ideological consistency [LC02]. Generally, the political elite are perceived as exhibiting greater ideological coherence and predictability in their positions across various domains compared to the mass. Some theories (e.g. [LC02, KK17]) suggest that the public are actually immune to elite polarization, whereas political elites in the opposing groups, are growing more ideologically distant from each other while also becoming more internally homogeneous [RM16]. However, recent body of work has found people’s views converging and becoming more interwined, possibly indicating that certain ideological realignment is taking place among the mass too [KM21]. To our knowledge, prior research on distinguishing the contribution of the elite and the mass to distinct aspects of polarization in online systems is rather limited. Our approach will help us evaluate whether polarization among the mass has remained at the same level, independent of the polarization trends observed among the elite.
Additionally, there exists work underscoring the importance of reconsidering the idea of treating both sides of a political discussion as equal [JBD22, GR14, SRZ19]. The language [KC24], rhetoric, and distribution of power can vary significantly between different ideological groups. Equivalently, polarization can be also asymmetrical [JBD22], which indicates that a group can become more extreme while the other group staying at the same level. The majority of current methodologies struggle to discern which group or groups contribute more significantly to overall structural polarization or issue alignment. Implementing a consistent partition strategy with an intuitive characterization of identified structures allows for a more precise description of disparities between these polarized groups and enables comparisons across different topics represented by networks.
Generally, assuming that each individual or group contributes uniformly to polarization can be suboptimal for several reasons. As the literature has shown, we should expect that the nature of an individual’s divisiveness depends on their role in society. We also know that the mass can take social and political cues from political elites [VBS21, DL19, GEN+20, Sky21], and are often influenced by them [All23, MS21, KT18, GEN+20, Ber07, MS18]. When elites indicate a divergence in ideology or display hostility towards out-groups, their followers interpret these signals and may adjust their behavior accordingly [BC18, FH23, Ber07, MS18]. This phenomenon, commonly referred to as elite signaling or elite influence, has been linked to political polarization [BCRR23, Sky21]. While the impactful role of elite in political and social systems is evident, one should not underestimate the impact of the mass either. Previous research suggests that politically-engaged individuals within the public may be the ones setting the political agendas for discussions on online platforms [BCN+19]. Not only might the people be leading the conversations, they are often needed for reaching wider audience in the social networks [BWB+15]. While the connection between these unique characteristics of the elite and mass groups and polarization is complex and, for the most part, unknown, it is crucial to examine how they differ in their contribution to the political polarization observed online.
In recent decades, three distinct aspects of polarization have received significant attention in the community: 1) the bimodality of opinion distributions, 2) the extent of issue and partisan alignment, and 3) the affective dimension, which refers to the development of strong negative attitudes and emotions towards members of opposing political or ideological groups. While these three aspects of polarization can potentially strengthen each other, it is important to recognize that they are different concepts, both theoretically and in terms of how they are measured empirically [HDC23]. In this paper, we focus on the first two aspects.
Researchers have traditionally examined the bimodality and alignment aspects of polarization using survey-based data, estimating ideological shifts and correlations between distinct issue positions from participants’ responses (e.g. [KM21]). However, the research community has increasingly turned its attention to social media platforms, as, in today’s world, these platforms serve as crucial venues and spaces for daily discussions on various political topics and debates. Additionally, these online platforms have become arenas for political competition, with millions of interactions and users, leading to valuable insights into the ideological distances among the users, their limited communication, and elite influence on public opinion worldwide (e.g. [Bri18, CRF+11, CMK+20, GEN+20]). Social media platforms enable researchers to monitor conversations and interactions in almost real-time, allowing for timely analysis of emerging trends and shifts in public opinion. Besides the inexpensive access to large datasets, it is also reasonable to expect that there are less response bias in users’ online behavior compared to traditional survey experiments. These advantages may explain the growing interest in analyzing polarization through the structural properties of networks.
Within social networks the bimodality of opinion distribution can be estimated by measuring the limited communication or reachability between the opposing groups in a given network [MBLB15], although some recent work have advocated for handling these separately [HDC23]. The intuition is that sharper and further peaks in an opinion distribution related to a topic X would correspond to a more isolated communities in a network where the topic X is being discussed. This type of isolation is often referred to as structural [SCK22] or interactional [FZQ+23] polarization. Most of the methods designed to capture this assesses the separation of interaction patterns among two distinct groups [GMGM18] — although some work have been recently done for handling multiple groups [NI22, MGLB23]. Typically, quantifying the structural polarization of a political communication network involves a three-step process [GMGM18, SCK22]. Firstly, the network is constructed based on the collected data. Following this, community detection techniques are employed to identify functional groups, often referred to as polarized groups, within the network, which frequently represent opposing sides of a political discourse. Finally, the quality of the division is evaluated using some structural polarization score, such as Random Walk Controversy [GMGM18] or Adaptive EI-index [CSG+21]. Again, the assumption these methods often have is that this division between the two clusters corresponds to either the lack of interactions between the clusters or the degree to which individuals are trapped within their own communities. Moreover, many of these methods require identifying the key figures in a network in order to quantify the extent of separation [GMGM18, SCK22, MBLB15], in which typically researchers rely on the number of endorsements a user has as a proxy for their importance or “eliteness”. While some polarization studies utilize known node labels [FGT+22, XGM+22] to identify e.g. politicians, this information is not always available and might miss other crucial users without clear affiliations, such as nonpolitical opinion leaders, which are shown to be more popular than the political figures on social media [MJL22]. Although our method is not specifically designed to address this potential limitation, there are situations where inferring the key figures in a network solely from certain structural properties may be preferable.
While these methods have produced valuable insights into societal division trends, their approach to measuring polarization may be somewhat restrictive. In addition to the inflexibility in choosing the number of groups, they are primarily designed to quantify the extent of interactional patterns within a single issue, and hence, not being capable of capturing the potential correlation between individuals’ stances on different topics. Some political and social polarization literature identifies issue alignment as more deleterious aspect of polarization [Mas15, CSG+21, Tör22]. Issue alignment refers to the degree of agreement or similarity between individuals or groups on a set of issues. In issue alignment, individuals or groups may align themselves based on shared values or ideologies leading to a collective stance on range of issues, potentially deepening divisions in the society [Mas15, Tör22].
Our contribution comprises both methodological and exploratory aspects. We study five topics that have become drastically more polarized online in Finland over the course of four years, spanning the period from 2019 to 2023, which contains major events, such as a global pandemic and a war in Europe. In this paper, we apply this framework to analyze polarization trends and highlight their differences between two snapshots of the Finnish Twittersphere. Specifically, we will examine the periods surrounding two parliamentary elections, in 2019 and 2023, and demonstrate how polarization dynamics vary across networks. We find that the polarized groups, as well as the elite and mass within them, shape polarization trends in very different ways.
To reiterate, our objective is to propose a method capable of measuring two distinct types of polarization—structural polarization and issue alignment—within both elite and public spheres, utilizing network representations of social or political systems. Additionally, our method unravels the imbalances in the impact that distinct polarized groups have on the observed polarization—something that a single polarization value would not reveal.
This paper is organized as follows: In Section 2, we outline the steps for inferring hierarchical structures in a polarized network and introduce our methods, which utilize these hierarchies to describe polarization. The main findings are discussed in Section 3, where we study the trends and patterns in hierarchical polarization on Twitter during the Finnish parliamentary elections in 2019 and 2023. Finally, in Section 4, we conclude with a discussion, list of identified limitations and ideas for future work.
2 Hierarchies in polarized networks
Figure 1: Observed network is depicted by the gray block, and it is divided into two assortative groups (pink and turquoise blocks) only when there is sufficient evidence of such groupings in the network. Once the polarized groups are identified, each block is further partitioned into a core and periphery independently, given significance of the hierarchy. Following the hierarchy decomposition, there are four potential scenarios: both polarized groups exhibit a hierarchical structure (1), one of the groups demonstrates a hierarchical structure (2-3), or neither of the groups displays indications of a hierarchical structure (4).
2.1 Finding the groups and hierarchies
@@
Polarization is often assumed to significantly impact social network structures, giving rise to two internally tightly-knit groups with sparse connections between them [GMGM18, HDC23]. Therefore, the level of polarization in a system can be quantified by examining the patterns of interactions within that configuration [FGT+22, GMGM18]. To conduct our experiments, our initial step is to identify the polarized groups. Following this, we need to distinguish between the elite and the mass members within each polarized group. This means that each individual in a network belongs to a specific polarized group (for-against) and a hierarchical group (elite-mass). While from now we will focus on two polarized groups for this study, this method can be readily adapted to scenarios involving more than two opposing groups.
Decomposing a network involves breaking it into modular and structurally distinct components. The conventional approach typically entails identifying the polarized groups in a network using clustering algorithms [GMGM18]. However, prior research has highlighted that many clustering algorithms can generate convincing group structures even in entirely random networks [LRR10, GSPA04, ZM14]. This tendency results in non-controversial networks having inflated polarization values, leading to a high number of false positives [SCK22]. To alleviate this issue, we employ the stochastic block model (SBM) to identify both polarized groups and hierarchical groups in a network, exploiting the clustering method’s ability to compute the description length for the given model and partition. Minimum description length is a principle for model selection that favors models with the shortest code length needed to describe the data, consistent with Occam’s razor, which suggests simpler explanations are generally preferable [Grü07]. Thus, an observed community structure that could be explained due to randomness, would be flagged with SBM, as the description length for an SBM with no community structure would be lower than that of a model attempting to overfit by dividing the network into two communities [Yan16, Pei19, Pei14b, ZP20]. Note, however, that this does not completely eliminate the need of normalizing the polarization scores for instance in classification tasks [SCK22].
There are two assumptions that are translated into our choices of SBM models. According to our first assumption a polarized network consists of two assortative groups, and thus, we apply the planted partition model variation of the SBM [ZP20, Pei14b]. We sweep the number of blocks (groups) between 1 to 2, and select the configuration leading to the lowest description length.
@@
Our second assumption is the existence of hierarchies in each polarized group. Social groups tend to naturally self-organize into hierarchies, where members exhibit varying levels of influence, expertise, or dominance [KXO15]. The emergence of these hierarchies is considered both inevitable and often advantageous for social cohesion [KXO15, MG08]. We utilize the well-established concept of core-periphery structures (see e.g. [RPFM17]) in approximating the hierarchical nature of these polarized groups. The core-periphery structure is a common pattern observed in various systems, including social networks [BE00, KM17, YZS+18], economics [CLWU13, Wan16], and many other fields [CLWU13, KM17]. It consists of a densely interconnected core, typically representing nodes with higher connectivity, influence, or resource access, surrounded by a sparsely connected periphery representing nodes with lesser influence. In our context, the core is equivalent to the elite, and the periphery is equivalent to the mass (or public).
@@
We anticipate that the elite are typically positioned within structurally dominant areas of networks, forming cohesive structures [MJRW20]. On Twitter, the elite often include influential figures such as politicians, journalists, activists, and experts who are active on social media. In contrast, we assume the surrounding periphery comprising the mass, consisting primarily of users who endorse or amplify the core’s voice [BWB+15]. These assumptions are later assesed for our data in Section 3.
@@
For identifying the hierarchical structures that are embedded in the already identified polarized groups, we follow the method proposed by Gallagher et al. to infer the core-periphery structure [GYW21] within the SBM framework. We treat each group independently, which enables us to compare their observed hierarchies to the null model separately. We use the common Erdős–Rényi model as the baseline, as done for instance in [KM18], because it assumes no inherent hierarchy, even one that might be explained by the network’s degree sequence. This approach has its pros and cons that we address separately in the Appendix A.7. A visual depiction of the overall network decomposition pipeline is shown in Figure 1. Note that alternative methods for identifying core-periphery structures exist, such as the extended version of the original method [BE00] for finding multiple core-periphery pairs [KM17] or the method based on detecting core-periphery structures by surprise [dJCS19].
(a) Full network
(b) Cores
(c) Peripheries
Figure 2: The complete retweet network representing Climate discussion (2023) in (a), subgraph containing the identified cores in (b), and subgraph containing peripheries in (c). The complete observed graph is drawn with ForceAtlas2 -layout algorithm [JVHB14]. For the others graphs, same node and edge positions are applied with specific hierarchy groups filtered out.
2.2 Defining structural polarization and issue alignment for hierarchies

### v2

@@ Abstract
In the political arena of social platforms, opposing factions of varying sizes show asymmetrical patterns, and elites and masses within these groups have divergent motivations and influence, challenging simplistic views of polarization. Yet, existing methods for quantifying polarization reduce division to a single value, assuming uniform distribution of polarization online. While this approach can confirm the observed increase in political polarization in many societies, it overlooks complexities that could explain this phenomenon. Notably, opposing groups can have unequal impacts on polarization, and the literature shows division between elites and the masses is a critical factor to consider.
We propose a method to decompose existing polarization measures in order to quantify the role of groups, determined by these distinct hierarchies, in the total polarization value. We applied this method to polarized topics in the Finnish Twittersphere surrounding the 2019 and 2023 parliamentary elections. Our analysis reveals two key insights: 1) The impact of opposing groups on observed polarization is rarely balanced, and 2) while elites strongly contribute to structural polarization and consistently display greater alignment across various topics, the masses have also recently experienced a surge in issue alignment, a stronger form of polarization.
Our findings suggest that the masses may not be as immune to an increasingly polarized environment as previously thought. This research provides a more nuanced understanding of polarization dynamics, offering potential insights into its underlying mechanisms and evolution.
@@ 1 Introduction
In an era of increasing political polarization [57, 28, 8, 31, 19, 30, 49], the ability to dissect and understand the intricate structures within social systems has never been more crucial. The proliferation of social media platforms has transformed the landscape of political discourse, creating vast digital arenas where polarization can be observed, measured, and analyzed in unprecedented detail.
Current approaches to measuring polarization often reduce complex social dynamics to a single dimension, failing to expose the hidden patterns behind the observed societal divisions. While these methods have been successfully applied to report polarization trends online [22], they fall short of capturing the full picture. This oversimplification obscures the multifaceted nature of polarization, where different groups and their members may contribute to and experience division in distinct ways.
This work proposes a method that utilizes existing polarization scores and network representations of social systems to measure two key types of polarization—structural polarization and issue alignment—separately for the elites and masses in polarized systems. Our approach not only unveils significant imbalances between polarized groups but also highlights the disproportionate role of elites in shaping polarized environments.
The distinction between elite and mass polarization is crucial for understanding the true nature of political division in modern democracies. While headlines and public discourse often paint a picture of a deeply divided populace, this perception may be skewed by the outsized influence of political elites in media narratives. By conflating elite and mass polarization, we risk misdiagnosing the health of our democratic systems and implementing misguided solutions. Separating these measurements allows us to discern whether the apparent chasm in political attitudes truly reflects widespread societal division or if it predominantly exists among a small, albeit influential, segment of the population.
In addition, the relationship between political elites and the public is characterized by substantial asymmetry, with elites often serving as influential cue-givers [39, 16, 23, 58, 61] who shape mass attitudes and behaviors [5, 39, 35, 43, 44, 23, 1]. This elite signaling phenomenon, where followers adjust their stances in response to elite ideological divergence or out-group hostility [5, 3, 43, 20], has been linked to political polarization [58, 2]. Traditionally, elites have been viewed as more ideologically coherent and predictable than the general public, with some theories (e.g. [38, 29]) even suggesting that masses remain resistant to elite polarization while opposing elite groups grow increasingly distant and internally homogeneous [54]. However, recent research complicates this narrative, revealing potential ideological realignment among the masses [39, 36]. Our method enables tracking of these complex dynamics on online platforms, offering a real-time lens into whether and how elite-mass polarization patterns are evolving in digital spaces.
This study was motivated by the rapid increase in political polarization observed in Finland over a relatively short period. We examine five topics that have become significantly more polarized online in Finland between 2019 and 2023, a period marked by major global events including a pandemic and a war in Europe. By analyzing polarization trends across two snapshots of the Finnish Twittersphere—centered around the parliamentary elections of 2019 and 2023—we reveal distinct polarization dynamics across different networks. Our findings demonstrate that polarized groups, and the elites and masses within them, shape polarization trends in markedly different ways. This analysis not only provides insight into Finland’s evolving political landscape but also offers a method for understanding polarization dynamics in other contexts.
The paper is structured as follows: Section 2 briefly reviews current polarization measures. Section 3 outlines our method for inferring hierarchical structures in polarized networks, and decomposes polarization using these hierarchies. Section 4 presents our main findings, analyzing hierarchical polarization trends on Twitter during the 2019 and 2023 Finnish parliamentary elections. We conclude with a discussion in Section 5.
2 Polarization Measures
Quantifying political polarization in society is a complex task. Traditional approaches have relied on survey-based data to measure opinion bimodalities and issue alignment, estimating ideological shifts and correlations between distinct issue positions (e.g., [36]). However, the research landscape has significantly evolved with the emergence of social media platforms, which now serve as critical arenas for daily political discourse and debate. This digital transformation has prompted researchers to increasingly focus on these platforms as rich sources of data for understanding contemporary political dynamics [9, 11, 12, 23].
In social networks, opinion distribution bimodality can be estimated by measuring communication limitations or reachability between opposing groups [45, 27]. The underlying assumption is that sharper, more distant peaks in opinion distribution correlate with more isolated communities in topic-specific discussion networks. This isolation is often termed structural [56] or interactional [19] polarization. Most methods designed to capture this phenomenon assess interaction pattern separation between two distinct groups [22], although recent work has expanded to handle multiple groups [48, 41].
Measuring structural polarization in political communication networks typically involves a three-step process: constructing a network from collected data, identifying functional groups using community detection techniques, and evaluating the division strength with structural polarization scores [22, 56]. Common metrics for this final step include Random Walk Controversy [22] and Adaptive EI-index [10], which assess the degree of inter-group isolation or intra-group cohesion. The common intuition these scores hold is to capture the extent to which individuals are “trapped” within their own ideological communities or avoid engaging with opposing viewpoints.
Many polarization quantification methods require identifying key network figures, often using endorsement counts (e.g., likes or reshares) as a proxy for importance or “eliteness” [22, 56, 45]. While some studies utilize known node labels to mark politicians in a network [18, 63], this approach may overlook crucial users without clear affiliations, such as nonpolitical opinion leaders, who often outpace political figures in social media popularity [47]. Although our method wasn’t specifically designed to address this limitation, inferring key figures solely from structural properties can be advantageous in certain scenarios, particularly when explicit labeling is unavailable or insufficient.
Structural polarization scores, while insightful, often have limitations: they typically focus on two-group systems and single-issue interaction patterns, overlooking the critical aspect of issue alignment in multi-group environments. Issue alignment—the degree of agreement across multiple topics—is increasingly recognized as a more pernicious form of polarization [42, 10, 59]. It occurs when individuals or groups adopt collective stances on a range of issues based on shared values or ideologies, potentially exacerbating societal divisions [42, 59]. Researchers have applied more classic statistical methods, such as correlation [36], and information theory-based scores for quantifying issue alignment in polarized systems [10].
3 Groups and Hierarchies in Polarized Networks
@@
Polarization is often assumed to significantly impact social network structures, giving rise to two internally tightly-knit groups with sparse connections between them [22, 27]. Measuring polarization separately for elites and the masses requires the inference of their group memberships (for-against) and hierarchy (elite-mass). While from now we will focus on two polarized groups for this study, this method can be readily adapted to scenarios involving more than two opposing groups or hierarchies.
We begin by presenting the steps for identifying polarized groups in a network. Next, we partition these groups into smaller hierarchical subgroups based on predefined connectivity patterns and conceptualize their possible interactions in a polarized environment. Finally, we connect the opposing groups, their hierarchies, and actions to existing polarization measures.
3.1 Identifying the Polarized Groups
The conventional approach for finding the assortative groups in polarized network is using clustering algorithms [22]. We employ the stochastic block model (SBM), which is a probabilistic model used to represent and analyze community structure in networks by grouping nodes into blocks, where the probability of connections between nodes depends solely on their block memberships. We specifically use the contrained version of this model, also known as planted-partition model [51, 67, 53], which has been previously applied to finding polarized groups [53].
The generative nature of SBM enables rigorous model selection, as one can apply the Occam’s razor principle to select the configuration with lowest description length [24]. The property is particularly important as an observed community structure that could be explained due to randomness [26], can be flagged, as the description length for an SBM with no community structure would be lower than that of a model attempting to overfit by dividing the network into two communities [51, 64, 52, 67]. Therefore, when combined with model selection, SBM-based approach would result in zero polarization in entirely random networks, which is in contrast to most commonly used polarization pipelines introduced in the literature [26, 56]. We sweep the number of blocks (groups) between 1 to 2, and select the configuration leading to the lowest description length.
3.2 Unraveling the Hierarchical Structure
@@
Our second assumption is the existence of hierarchies in each polarized group. Social groups tend to naturally self-organize into hierarchies, where members exhibit varying levels of influence, expertise, or dominance [34, 60]. The emergence of these hierarchies is considered both inevitable and often advantageous for social cohesion [34, 40].
@@
We utilize the well-established concept of core-periphery structures (see e.g. [55]) in approximating the hierarchical nature of these polarized groups. The core-periphery structure is a common pattern observed in various systems, including social networks [7, 32, 66], economics [13, 62], and many other fields [13, 32]. It consists of a densely interconnected core, typically representing nodes with higher connectivity, influence, or resource access, surrounded by a sparsely connected periphery representing nodes with lesser influence. In our context, the core can be interpreted as representing the elites, with the periphery acting as the masses. This is in align with the elites typically positioned within structurally dominant areas of networks, forming cohesive structures [46]. On Twitter, elites often include influential figures such as politicians, journalists, activists, and experts who are active on social media. In contrast, we assume the surrounding periphery comprising the mass, consisting primarily of users who endorse or amplify the core’s voices [4]. These assumptions are later assessed for our data in Section 4.
@@
For identifying the hierarchical structures that are embedded in the already identified polarized groups, we follow the method proposed by [21] to infer the core-periphery structure [21] within the SBM framework. We treat each group independently, which enables us to compare their observed hierarchies to the null model separately. We use the common Erdős–Rényi model as the baseline, as done for instance in [33], because it assumes no inherent hierarchy, even one that might be explained by the network’s degree sequence. This approach has its pros and cons that we address separately in the Appendix B. Note that alternative methods for identifying core-periphery structures exist, such as the extended version of the original method [7] for finding multiple core-periphery pairs [32] or the method based on detecting core-periphery structures by surprise [14].
3.3 Interactions in Polarized Environments
Figure 1: Our conceptualization enables categorizing the actions that lead to an increase in the structural polarization score. (A) is an example of a polarized network that has been partitioned into two groups representing communities with distinct stances on a specific political topic. Large nodes correspond to the elite members, while smaller ones indicate the mass. In (B), a new connection is formed between two elite members, making the elite group more cohesive. Another polarization increasing action is when a new connection is formed towards the elite group either from an existing node or a new node. This is depicted in (C) and the overall degree of this type of actions is called mass amplification. Lastly, a connection between two members belonging to the mass, is shown in (D), illustrating the interpretation of mass cohesion. Lower mass cohesion can be seen corresponding to higher centralized opinion leadership among the elites, as most connections are directed towards them.
A common approach is to use a single score to measure the degree of division between these polarized groups (Fig. 1A). Now, instead of treating all individuals uniformly, we label the strongly internally connected nodes as elites (larger circles), and nodes that have few or weak connections among themselves but well-connected to the central of their group are labeled as the masses (smaller circles). This conceptualization enables us categorize the following interactions:
Elite-Elite: Each new connection between two elites within own group can be deemed as turning the core more connected, and therefore, increasing the elite cohesion of that group (Fig. 1B).
Mass-Elite: Each elite member can receive an endorsement or validation from the surrounding mass, increasing the total mass amplification of the corresponding group in the system (Fig. 1C).
Mass-Mass: A polarized group may have a very centralized opinion leadership, where mostly elites voices are amplified or more heterogeneous amplification patterns depending on the group’s mass cohesion (Fig. 1D), which increases as a connection is formed between two mass members.
3.4 Defining Structural Polarization and Issue Alignment for Groups and Hierarchies